# User Feedback Alignment for LLM-powered Exploration in Large-scale Recommendation Systems

### v3

@@ 4.2 Inference Scaling with Best-of-N User Alignment
We use the user alignment model as a user surrogate to critique the relevancy of the novel clusters predicted by the novelty LLM, which is lightly finetuned with users’ interaction histories. To increase the chance of predicting a novel cluster that is more aligned with user preference, we repeatedly and independently sample 5 times more predictions from the novelty LLM with high temperature, and then rank the predictions using the alignment model and pick the top k where k is the number of clusters served by the production system. Because the novelty LLM sampling, reward model scoring, and the best-of-n selection all happens offline, and we serve the same number of clusters in live traffic, there is no latency impact, and the additional cost of scaling up inference is amortized across offline bulk inference runs.
Getting a guarantee of the novelty of the prediction is important for user interest exploration. The repeated sampling of the novelty LLM improves the reasoning quality and maintains the prediction novelty while the alignment model selects the predictions users may prefer. This dual LLM setup avoids the challenge of teaching LLM both novelty and relevancy, which are two competing objectives and may result in catastrophic forget. By reflecting on the novelty prediction using an LLM aligned with user feedback, we improve the exploration efficiency by demoting the predictions that may result in lower user satisfaction.
@@ 5.1 Experimental Setup
Baseline. Besides comparing to the baseline novelty model without user alignment, we also compare the proposed method to existing production models: (1) Exploration-oriented models include: Hierarchical contextual bandit Song et al. (2022) obtain the next clusters through a tree-based LinUCB; Neural linear bandit-based DNN model Su et al. (2024) to predict the next novel cluster. Although these models are tailored to explore user interests, they are trained on interest transitions existing in the system and therefore are still subject to the feedback loop. (2) Exploitation-oriented models include a regular two-tower model Yang et al. (2020) and transformer-based Chen et al. (2019); Shaw et al. (2018) sequential model trained on all positive user feedback. Our live experimental results demonstrate our proposed method can lead to recommendation which are more novel and of better quality compared to these existing models.
@@ 5.2 Model Finetuning and Offline Evaluation
We used offline metrics to guide model training, checkpoint selection, and hyper-parameter searching (e.g. score normalization strategy). Offline evaluation is done on a holdout set of interest cluster sequences, the novel interest transitions and user’s feedback scores. During evaluation, the alignment reward model scores and ranks the interest cluster transitions for each input cluster sequence. We compare this model-generated ranking against the ground-truth ranking from the user feedback. Performance is measured using F1@K (i.e., the harmonic mean of precision and recall), and NDCG@K metrics, with K being the number of interest clusters served in live traffic.
As shown in Figure 3, the offline metrics improve consistently over a random baseline throughout the alignment model’s training process. These results underscore the importance of incorporating user feedback alignment into our inference scaling approach. Furthermore, the offline evaluation guided the hyper-parameter tuning, allowing us to optimize the reward model’s performance and prevent overfitting. In live A/B experiments, we deployed two arms: one favorable arm using an alignment model trained for 50,000 steps (where F1 converged in offline evaluation as shown in Figure 3), and another arm using an alignment model trained for 100,000 steps beyond the favorable converging point as comparison. We observed significantly improved user satisfaction with the favorable arm, as evidenced by a larger positive playback rate (PPB) gain – indicating better alignment with user preferences. This finding is consistent with our hypothesis that extensive training beyond the convergence point can lead to overfitting. While NDCG encourages the model to reproduce the exact ranking from user feedback, F1@K focuses on the model’s ability to identify the top-K most relevant clusters, which is more crucial for our top-n selection task. Memorizing the exact rankings is unnecessary and potentially detrimental to the exploration of novel and engaging recommendations.
@@ 5.3 Results and Analysis
Novelty and Quality. In Figure 4 (a), we compare the proposed method with various baseline models currently in production. Using the performance of Hierarchical contextual bandit Song et al. (2022) as the base, we measure improvement of novelty and quality of other models in our system. Specifically, we plot the increase in the novel impression ratio (impressions from interest clusters the user has never interacted with) to highlight recommendation novelty (x-axis), and the increase in positive playback rate to demonstrate recommendation quality (y-axis). We observed that aligning the novelty-focused novelty model with user preference results in higher users’ positive playback ratio at a slight cost of novelty. Nonetheless, the proposed method still has the highest novel impression ratio compared to the rest of the system. Additionally, our method achieves significantly better quality than existing exploration-oriented methods, even surpassing the exploitation-oriented methods. It is rare in the recommendation systems to achieve high novelty and user satisfaction simultaneously. This means through user feedback alignment, we moved our model to a more optimal point in the operation curve – over user satisfaction and engagement improved while the novelty is still the highest in the system.
@@
User Interest Exploration. To measure if the recommenders encourage users to explore new interests, we use unique engaged user-cluster (UEUC), which tracks the number of unique user-cluster engagement pairs. Figure 4(b) shows that our proposed user feedback alignment method not only improves the user satisfaction but also improves the number of user interests. This means our method improves the exploration efficiency. We also notice UEUC is higher for more active users, potentially because the reward model aligns more closely with the preferences of core users who contribute a larger portion of the user feedback training data.
@@ 6 Conclusion
In this paper, we leverage the hierarchical planning paradigm of LLMs for large-scale recommender systems, separating the exploration and exploitation goals into two distinct models during high-level planning. We share our successful approach to improving alignment using collective user feedback gathered from LLM-powered recommendation systems. Live experiments on a large-scale recommendation platform demonstrate that our proposed method enhances exploration efficiency while simultaneously increasing user engagement.

### v4

@@ 4.2 Inference Scaling with Best-of-N User Alignment
We use the user alignment model as surrogate for user preference to critique the relevancy of the novel clusters predicted by the novelty LLM, which itself is lightly fine-tuned with users’ interaction histories. To increase the chance of predicting a novel cluster that is more aligned with user preference, we repeatedly and independently sample 5 times more predictions from the novelty LLM with high temperature, and then rank the predictions using the alignment model and pick the top k where k is the number of clusters served by the production system. Because the novelty LLM sampling, reward model scoring, and the best-of-n selection all happens offline, and we serve the same number of clusters in live traffic, there is no latency impact, and the additional cost of scaling up inference is amortized across offline bulk inference runs.
Maintaining the novelty of predictions is crucial for effective user interest exploration. The repeated sampling of the novelty LLM improves the reasoning quality and maintains the prediction novelty while the alignment model selects the predictions users may prefer. This dual LLM setup avoids the challenge of teaching an LLM both novelty and relevancy – two competing objectives that can risk catastrophic forgetting. By evaluating the novelty prediction using an LLM aligned with user feedback, we improve the exploration efficiency by demoting the predictions that may result in lower user satisfaction.
@@ 5.1 Experimental Setup
Baseline. Besides comparing to the baseline novelty model without user alignmentWang et al. (2024c), we also compare the proposed method to existing production models: (1) Exploration-oriented models include: Hierarchical contextual bandit Song et al. (2022) obtain the next clusters through a tree-based LinUCB; Neural linear bandit-based DNN model Su et al. (2024) to predict the next novel cluster. Although these models are tailored to explore user interests, they are trained on interest transitions existing in the system and therefore are still subject to the feedback loop. (2) Exploitation-oriented models include a regular two-tower model Yang et al. (2020) and transformer-based Chen et al. (2019); Shaw et al. (2018) sequential model trained on all positive user feedback. Our live experimental results demonstrate our proposed method can lead to recommendations that are more novel and of better quality compared to these existing models.
@@ 5.2 Model Finetuning and Offline Evaluation
We used offline metrics to guide the alignment model training, checkpoint selection, and hyper-parameter searching (e.g. score normalization strategy). Offline evaluation is done on a holdout set of interest cluster sequences, the novel interest transitions and user’s feedback scores. During evaluation, the alignment reward model scores and ranks the interest cluster transitions for each input cluster sequence. We compare this model-generated ranking against the ground-truth ranking from the user feedback. Performance is measured using F1@K (i.e., the harmonic mean of precision and recall), and NDCG@K metrics, with K being the number of interest clusters served in live traffic.
As shown in Figure 3, the offline metrics improve consistently over a random baseline throughout the alignment model’s training process. These results underscore the importance of incorporating user feedback alignment into our inference scaling approach. Furthermore, the offline evaluation guided the hyper-parameter tuning, allowing us to optimize the reward model’s performance and prevent overfitting. In live A/B experiments, we deployed two arms: one favorable arm using an alignment model trained for 50,000 steps (where F1 converged in offline evaluation as shown in Figure 3), and another arm using an alignment model trained for 100,000 steps beyond the favorable converging point as comparison. We observed significantly improved user satisfaction with the favorable arm, as evidenced by a larger positive playback rate gain – indicating better alignment with user preferences. This finding is consistent with our hypothesis that extensive training beyond the convergence point can lead to overfitting. While NDCG encourages the model to reproduce the exact ranking from user feedback, F1@K focuses on the model’s ability to identify the top-K most relevant clusters, which is more crucial for our top-n selection task. Memorizing the exact rankings is unnecessary and potentially detrimental to the exploration of novel and engaging recommendations.
@@ 5.3 Results and Analysis
This section shows our method simultaneously improves recommendation novelty and user satisfaction, outperforming baselines in engagement and exploration, and details the benefits of its production-deployed pointwise labeling strategy.
Novelty and Quality. In Figure 4 (a), we compare the proposed method with various baseline models currently in production. Using the performance of Hierarchical contextual bandit Song et al. (2022) as the base, we measure improvement of novelty and quality of other models in our system. Specifically, we plot the increase in the novel impression ratio (impressions from interest clusters the user has never interacted with) to highlight recommendation novelty (x-axis), and the increase in positive playback rate to demonstrate recommendation quality (y-axis). We observed that aligning the novelty model with user preference results in higher users’ positive playback ratio at a slight cost of novelty. Nonetheless, the proposed method still has the highest novel impression ratio compared to the rest of the system. Additionally, our method achieves significantly better quality than existing exploration-oriented methods, even surpassing the exploitation-oriented methods. It is rare in recommendation systems to achieve high novelty and user satisfaction simultaneously. This means through user feedback alignment, we moved our model to a more optimal point in the operation curve – over user satisfaction and engagement improved while the novelty is still the highest in the system.
@@
User Interest Exploration. To measure if the recommender encourage users to explore new interests, we use unique engaged user-cluster (UEUC), which tracks the number of unique user-cluster engagement pairs. Figure 4(b) shows that our proposed user feedback alignment method not only improves the user satisfaction but also improves the number of user interests. This means our method improves the exploration efficiency. We also observed UEUC is higher for more active users, potentially because the reward model aligns more closely with the preferences of core users who contribute a larger portion of the user feedback training data.
@@ 6 Conclusion
In this paper, we advanced the hierarchical planning paradigm for LLM-powered large-scale recommendation systems by decoupling high-level planning into two specialized models: one focused on generating novel interest candidates and another focused on aligning these candidates with user feedback. We share our successful approach to improving alignment using collective user feedback gathered from LLM-powered recommendation systems. Live experiments on a large-scale recommendation platform demonstrate that our proposed method enhances exploration efficiency while simultaneously increasing user engagement.